OpenStudy (anonymous):

What is the solution of n^2- 49 = 0 A) -7 B) 7 c) +/- 7 D) no solution

OpenStudy (anonymous):

@mathman101

OpenStudy (anonymous):

@the_ocean_girl

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

; i have no idea tbh ,

OpenStudy (anonymous):

well i'd say c

OpenStudy (anonymous):

; why c ?

OpenStudy (anonymous):

you can solve using the quadratic formula

OpenStudy (anonymous):

X^2 = 64 = 0

OpenStudy (anonymous):

you want me to solve that one too

OpenStudy (anonymous):

; yes

OpenStudy (anonymous):

you can add 49 to both sides like this.. n² – 49 +49= 0+49

OpenStudy (anonymous):

n=7 and n=-7

OpenStudy (anonymous):

what are the answer choices

OpenStudy (anonymous):

@myalina

OpenStudy (anonymous):

A) -5 B) 8 C) +/- D) no solution @mathman101

OpenStudy (anonymous):

im not sure what do you think

OpenStudy (trojanpoem):

I agree with @mathman101 , +- 7

OpenStudy (anonymous):

i think b what do you think @TrojanPoem

OpenStudy (trojanpoem):

I think +-7 even if we plugged both (-7)^2 = 49 (7)^2 = 49, why do you think it's b ?

OpenStudy (anonymous):

no i mean the next question

OpenStudy (trojanpoem):

where is that ?

OpenStudy (anonymous):

X^2-64=0

OpenStudy (anonymous):

A) -5 B) 8 C) +/- D) no solution

OpenStudy (trojanpoem):

X^2 = 64 x = _+ 8 , c

OpenStudy (anonymous):

sooo you agree?

OpenStudy (trojanpoem):

I agree it's C

OpenStudy (anonymous):

oh you think c

OpenStudy (trojanpoem):

Why do you think it's b ?

OpenStudy (anonymous):

sorry didn't know the answer just guessed

OpenStudy (anonymous):

but now i see its c

OpenStudy (anonymous):

well i have to go good luck @myalina

OpenStudy (anonymous):

Its B ive done this b4

OpenStudy (anonymous):

3x^2 = 0 A) X = +/-3 B) X = +/- 1/3 C) X = 0 D) no solution

OpenStudy (anonymous):

@MIles3636

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@myalina u need anymore help

OpenStudy (trojanpoem):

3x^2 = 0 , x^2 = 0 , x = 0 (C)

OpenStudy (anonymous):

one of yal are wrong .. i dont wanna get the wrong answer @TrojanPoem @MIles3636

OpenStudy (anonymous):

Hes Wrong I did My math

OpenStudy (trojanpoem):

@MIles3636 , None of us is wrong, you see the zero as one 3x^2 = 0

OpenStudy (trojanpoem):

@myalina ,it's 3x^2 = 1 or zero ?

OpenStudy (anonymous):

1/4x^2 - 4 = 0 A) X = +/-3 B) X = +/-4 C) X = +/-2 D) no solution @MIles3636

OpenStudy (anonymous):

it says 0 @TrojanPoem

OpenStudy (trojanpoem):

So the answer is C, he has just seen the zero as one which resulted in miscalculating.